# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 4.1. $L^{2}$-uniform bound of the logarithmic derivative $\mathsf{d}^{N}$ of $\mu^{N}$
Note that $\mathfrak{m}_{R,k,l}^{N,[1]}({x},\mathsf{s})$ is symmetric in $(x,s^{1},\ldots,s^{k})$ when regarded as a function on $\overline{S}_{R}^{k+1}$ by the identification $({x},\mathsf{s})=({x},s^{1},\ldots,s^{k})$, where $\mathsf{s}=\sum_{i=1}^{k}\delta_{s^{i}}$. Let
$\mathbb{D}_{R}^{[1]}[\mathsf{h}]({x},\mathsf{s})=\frac{1}{2}\Big{(}\sum_{{x}\in S_{R}}|\partial_{{x}}\mathsf{h}({x},\mathsf{s})|^{2}+\sum_{s^{i}\in S_{R}}|\partial_{s^{i}}\mathsf{h}({x},\mathsf{s})|^{2}\Big{)}.$
Then, using the permutation invariance of $\mathfrak{m}_{R,k,l}^{N,[1]}$ in $(x,s^{1},\ldots,s^{N-1})$ and noting that $k+1$ particles exist in $\overline{S}_{R}$ for $({x},\mathsf{s})\in\overline{\mathsf{S}}_{R,k}^{[1]}$, we obtain
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathbb{D}_{R}^{[1]}[\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N,[1]}]d\Lambda_{R}^{[1]}\leq 2\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathbb{D}_{R}^{[1]}[\mathsf{f}_{Q}]\big{(}\mathfrak{m}_{R,k,l}^{N,[1]}\big{)}^{2}+\mathsf{f}_{Q}^{2}\mathbb{D}_{R}^{[1]}[\mathfrak{m}_{R,k,l}^{N,[1]}]d\Lambda_{R}^{[1]}$
$=\int_{\mathbb{R}^{d}\times\mathsf{S}}\lvert\nabla\mathsf{f}_{Q}\rvert^{2}\big{(}\mathfrak{m}_{R,k,l}^{N,[1]}\big{)}^{2}d\Lambda_{R}^{[1]}+2\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}^{2}\mathbb{D}_{R}^{[1]}[\log\mathfrak{m}_{R,k,l}^{N,[1]}]\big{(}\mathfrak{m}_{R,k,l}^{N,[1]}\big{)}^{2}d\Lambda_{R}^{[1]}$
$=\int_{\mathbb{R}^{d}\times\mathsf{S}}\lvert\nabla\mathsf{f}_{Q}\rvert^{2}\big{(}\mathfrak{m}_{R,k,l}^{N,[1]}\big{)}^{2}d\Lambda_{R}^{[1]}+\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}^{2}(k+1)|\mathsf{d}^{N}|^{2}\big{(}\mathfrak{m}_{R,k,l}^{N,[1]}\big{)}^{2}d\Lambda_{R}^{[1]}$
@@
Hence, combining this with (4.5) and Lemma 4.3 yields the following.
$\limsup_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathbb{D}_{R}^{[1]}[\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}]d\Lambda_{R}^{[1]}<\infty.$ (4.19)
As before, we regard $\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}({x},\mathsf{s})$ as a function on $S_{R}^{k+1}$. Note that $\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}=0$ on $(\partial S_{R})\times S_{R}^{k}$ from (4.5) and $R\geq Q+1$. Hence, from the Rellich embedding theorem and (4.19), there exists $q>1$ such that $\{\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ is relatively compact in $L^{q}(S_{R}^{k+1},d{x}^{1}\cdots d{x}^{k+1})$, where we take 1<q<n/(n1)1<q<n/(n-1) for $n=d(k+1)$. According to this and (4.5), $\{\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ is relatively compact in $L^{q}(\Lambda_{R}^{[1]})$. From (4.16), $\{\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ is bounded in $L^{\infty}(\Lambda_{R}^{[1]})$. Hence, $\{\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ is relatively compact in $L^{2}(\Lambda_{R}^{[1]})$. This implies i.
From i and a diagonal argument, there exist a subsequence of $\{N_{\mathfrak{n}}\}$, denoted by the same symbol, and a limit $\{\widetilde{\mathfrak{m}}_{Q,R,k,l}^{[1]}\}$ such that
$\lim_{\mathfrak{n}\to\infty}\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}=\widetilde{\mathfrak{m}}_{Q,R,k,l}^{[1]}$ in $L^{2}(\Lambda_{R}^{[1]})$ (4.20)
for all $R\geq Q+1$ and $k\in\{0\}\cup\mathbb{N}$ and $l\in\mathbb{N}$. Then, we obtain the uniqueness of limit points of $\{\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ from the weak convergence $\lim\mu^{N_{\mathfrak{n}}}=\mu$ given in ((A2)). Hence, the full sequence converges to $\widetilde{\mathfrak{m}}_{Q,R,k,l}^{[1]}$ in (4.20). According to (4.5), we set
@@
From (4.20) and (4.21), we have (4.17). From (4.16) and (4.17), we obtain (4.18). ∎
Theorem 4.1.
$i$ $\mu$ has a one-point correlation function $\rho^{1}$.
@@
Using the consistency (4.18), we define the function $\mathfrak{m}_{R,k}^{[1]}$ by $\mathfrak{m}_{R,k}^{[1]}\wedge l=\mathfrak{m}_{R,k,l}^{[1]}$. Then, we define function $\mathfrak{m}_{R}^{[1]}$ by $\mathfrak{m}_{R,k}^{[1]}=1_{\overline{\mathsf{S}}_{R,k}^{[1]}}\mathfrak{m}_{R}^{[1]}$. Here, we used $\overline{\mathsf{S}}_{R,k}^{[1]}$ are disjoint sets. We define $\mu_{R}^{[1]}=\mathfrak{m}_{R}^{[1]}d\Lambda_{R}^{[1]}$. Then, $\mu_{R}^{[1]}$ becomes the one-reduced Campbell measure of $\mu_{R}$. Note that $\mathfrak{m}_{R,k}^{[1]}({x},s^{1},\ldots,s^{k-1}):=\mathfrak{m}_{R,k}^{[1]}({x},\sum_{i=1}^{k-1}\delta_{s^{i}})$ is symmetric on $\overline{S}_{R}^{k}$ for each $k\in\mathbb{N}$. Hence, we can construct the $k$-density function $\mathfrak{m}_{R,k}$ of $\mu$ on $\overline{S}_{R}$ from $\mathfrak{m}_{R}^{[1]}$ by the formula
@@
The constant $c_{4.4}$ is determined by the normalization due to the value $\mu_{R}(\overline{\mathsf{S}}_{R}^{k})$.
The one-point correlation function $\rho^{1}$ is given by the formula
$\rho^{1}({x})=\mathfrak{m}_{R,1}({x})+\sum_{j=2}^{\infty}\frac{1}{(j-k)!}\int_{\overline{S}_{R}^{j-1}}\mathfrak{m}_{R,j}({x},{x}^{2},\ldots,{x}^{j})d{x}^{2}\cdots d{x}^{j}.$ (4.22)
From (2.22), $\int_{\mathsf{S}}\mathsf{s}(S_{R})\mu(d\mathsf{s})<\infty$. Hence, the right-hand side of (4.22) is finite. Hence, $\rho^{1}$ is the one-point correlation function of $\mu$. ∎
4.4. The logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$.
This subsection aims to construct the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$. For this purpose, we begin by constructing the logarithmic derivative $\mathsf{d}_{R}$ of $\mu_{R}$, where $\mu_{R}=\mu\circ\pi_{\overline{S}_{R}}^{-1}$.
Let $\overline{\mathsf{S}}_{R,k}^{[1]}$, $\mathfrak{m}_{R}^{N,[1]}$, and $\mathfrak{m}_{R,k,l}^{N,[1]}$ be as in (4.16). Let
$\mathsf{M}_{R,k,l}^{N,[1]}$ $=\{({x},\mathsf{s})\in\overline{\mathsf{S}}_{R,k}^{[1]};0<\mathfrak{m}_{R}^{N,[1]}({x},\mathsf{s})<l\},$
$\mathsf{d}_{R,k,l}^{N}({x},\mathsf{s})$ $=1_{\mathsf{M}_{R,k,l}^{N,[1]}}({x},\mathsf{s})\nabla_{{x}}\log\mathfrak{m}_{R,k,l}^{N,[1]}({x},\mathsf{s}).$ (4.23)
Let $\mathfrak{m}_{R,k,l}^{[1]}$ be as in Lemma 4.6. Then, using (4.18), we define $\mathsf{M}_{R,k,l}^{[1]}$ and $\mathsf{d}_{R,k,l}$ by
$\mathsf{M}_{R,k,l}^{[1]}$ $=\{({x},\mathsf{s})\in\overline{\mathsf{S}}_{R,k}^{[1]};0<\mathfrak{m}_{R,k,l}^{[1]}({x},\mathsf{s})<l\},$
$\mathsf{d}_{R,k,l}({x},\mathsf{s})$ $=1_{\mathsf{M}_{R,k,l}^{[1]}}({x},\mathsf{s})\nabla_{{x}}\log\mathfrak{m}_{R,k,l}^{[1]}({x},\mathsf{s}).$ (4.24)
Lemma 4.7.
Let $\mathsf{f}_{Q}$ be as in (4.5). Then, for each $k\in\{0\}\cup\mathbb{N}$ and $R,l\in\mathbb{N}$,
$\lim_{\mathfrak{n}\to\infty}\mathsf{d}_{R,k,l}^{N_{\mathfrak{n}}}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]})^{1/2}$ $=\mathsf{d}_{R,k,l}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]})^{1/2}\quad\text{ weakly in $L^{2}(\Lambda_{R}^{[1]})$},$ (4.25)
Proof.
From Lemma 4.5, (4.16), and (4.23), we easily deduce that $\{\mathsf{d}_{R,k,l}^{N_{\mathfrak{n}}}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]})^{1/2}\}_{\mathfrak{n}}$ is bounded in $L^{2}(\Lambda_{R}^{[1]})$. Hence, we have a subsequence of $\{N_{\mathfrak{n}}\}$, denoted by the same symbol, and a limit $\widehat{\mathsf{d}}_{R,k,l}$ such that
From Lemma 4.6, we deduce
$(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]})^{1/2}=\lim_{\mathfrak{n}\to\infty}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]})^{1/2}$ $\text{ in }L^{4}(\Lambda_{R}^{[1]}).$ (4.28)
$\int_{S_{R}\times\mathsf{S}}\widehat{\mathsf{d}}_{R,k,l}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]})^{1/2}\mathsf{h}d\Lambda_{R}$
$=\lim_{\mathfrak{n}\to\infty}\int_{S_{R}\times\mathsf{S}}\mathsf{d}_{R,k,l}^{N_{\mathfrak{n}}}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]})^{1/2}(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]})^{1/2}\mathsf{h}d\Lambda_{R}.$ (4.29)
We set
$\widetilde{\mathsf{d}}_{R,k,l}({x},\mathsf{s})=\begin{cases}\widehat{\mathsf{d}}_{R,k,l}({x},\mathsf{s})(\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]}({x},\mathsf{s}))^{-1/2}&\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]}({x},\mathsf{s})>0\\ 0&\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]}({x},\mathsf{s})=0\end{cases}.$ (4.30)
Then, from (4.24) and (4.28)–(4.30), we obtain
@@
Hence, we have $\widetilde{\mathsf{d}}_{R,k,l}=\mathsf{d}_{R,k,l}$. From this, (4.27), (4.28), and (4.30), we obtain (4.25).
From (4.23), $\mathsf{d}_{R,k,l}^{N}$ satisfy $\mathsf{d}_{R,k,l}^{N}({x},\mathsf{s})=\mathsf{d}_{R,k,l+1}^{N}({x},\mathsf{s})$ for $({x},\mathsf{s})\in\mathsf{M}_{R,k,l}^{N,[1]}$ for all $l\in\mathbb{N}$. This property is inherited by $\mathsf{d}_{R,k,l}$ from (4.27)–(4.29). Hence, we obtain (4.26). ∎
Let $\mathsf{M}_{R}^{[1]}=\{({x},\mathsf{s});0<\mathfrak{m}_{R}^{[1]}({x},\mathsf{s})<\infty\}$ and $\mathsf{d}_{R}({x},\mathsf{s})=1_{\mathsf{M}_{R}^{[1]}}({x},\mathsf{s})\nabla_{{x}}\log\mathfrak{m}_{R}^{[1]}({x},\mathsf{s})$. Then, from (4.24), we derive the following:
@@
Lemma 4.8.
For all $R\geq Q+1$,
@@
$\lim_{\mathfrak{n}\to\infty}\mathsf{d}_{R}^{N_{\mathfrak{n}}}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}=\mathsf{d}_{R}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]})^{1/2}\quad\text{ weakly in $L^{2}(\Lambda_{R}^{[1]})$},$ (4.33)
$\lim_{\mathfrak{n}\to\infty}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}=(\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]})^{1/2}\quad\text{ in $L^{2}(\Lambda_{R}^{[1]})$}.$ (4.34)
Proof.
Recall that $\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]}d\Lambda_{R}^{[1]}=d\mu_{R}^{N_{\mathfrak{n}},[1]}$. Then, from Lemma 4.5, we obtain (4.32).
From (4.32), $\{\mathsf{d}_{R}^{N_{\mathfrak{n}}}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}\}_{\mathfrak{n}\in\mathbb{N}}$ is relatively compact under the weak convergence in $L^{2}(\Lambda_{R}^{[1]})$. From Lemma 4.7 and (4.31), the limit points are unique and coincide with $\mathsf{d}_{R}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]})^{1/2}$. Hence, we obtain (4.33).
From ((A2)), we have $\lim_{\mathfrak{n}\to\infty}\mu^{N_{\mathfrak{n}}}=\mu$ weakly. Hence, we obtain
$\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]}d\Lambda_{R}^{[1]}=\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]}d\Lambda_{R}^{[1]}<\infty.$
Then, from Lemma 4.6,
$\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\lvert(\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}-(\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]})^{1/2}\rvert^{2}d\Lambda_{R}^{[1]}$
$=$ $\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]}+\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]}-2(\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{[1]})^{1/2}d\Lambda_{R}^{[1]}=0.$
This implies (4.34). ∎
Let $\mathscr{F}_{R}^{[1]}=\mathscr{B}(\overline{S}_{R})\times\sigma[\pi_{\overline{S}_{R}}]$. Then, $\mathsf{d}_{R}$ is $\mathscr{F}_{R}^{[1]}$-measurable.
Theorem 4.2.
@@
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}d\mu_{R}^{[1]}=\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}^{N_{\mathfrak{n}}}d\mu_{R}^{N_{\mathfrak{n}},[1]},$ (4.36)
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}d\mu_{R}^{[1]}=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}\mathsf{h}d\mu_{R}^{[1]}$ (4.37)
for all $\mathscr{F}_{R}^{[1]}$-measurable $\mathsf{h}=(\mathsf{f}_{Q}\hat{\mathsf{f}})\otimes\mathsf{g}\in\mathscr{D}_{\circ\mathrm{b}}^{[1]}$ such that $\hat{\mathsf{f}}\in C_{0}^{\infty}(S_{R})$.
$i$ $\mathsf{d}_{R}$ is the logarithmic derivative of $\mu_{R}=\mu\circ\pi_{\overline{S}_{R}}^{-1}$.
@@
From Lemma 4.6, $d\mu_{R}^{[1]}=\mathfrak{m}_{R}^{[1]}d\Lambda_{R}^{[1]}$. Then, from (4.32) and Lemma 4.8,

### v4

@@ 5.2. Explicit formulae of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}$ of $\mu_{R,\mathsf{y}}$
$\int_{\mathsf{S}}\sum_{s^{i}\notin\overline{S}_{R+\epsilon}}^{\infty}\Big(\sup_{{x}\in S_{R}}\frac{1}{\lvert{x}-s^{i}\rvert^{d+\ell}}\Big)\,d\mu(\mathsf{s})$
@@
[…]
which is finite by (2.13). This together with $\Psi^{N}=\Psi$ and (1.1) proves (2.17). ∎
6. Non-collision in ISDEs with infinitely many particles
In this section, we prove the non-collision property of the stochastic dynamics for infinite-particle systems described by solutions of the ISDE (2.33). Importantly, we do not rely on properties of the associated quasi-regular Dirichlet forms, such as capacity.
The main result of this section, Proposition 6.1, applies to weak solutions of the ISDE (2.33). These solutions are independent of Dirichlet forms. We apply Proposition 6.1 to the solutions $\underline{\mathbf{X}}$ and $\underline{\mathbf{X}}_{\mathsf{a}}$ constructed in Subsection 8.2 and Subsection 10.1 via lower Dirichlet forms.
All results in Section 6 are proved under ((A1))((A3)).
Proposition 6.1.
Let $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ be a weak solution of (2.33). We define $\mathscr{X}_{u}^{i}=\big(X_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{X_{u}^{j}}\big)$. Assume that
$E\Big[\int_{0}^{T}\sum_{i\in\mathbb{N}}1_{S_{R}}(X_{u}^{i})\lvert\mathfrak{b}(\mathscr{X}_{u}^{i})\rvert^{2}du\Big]<\infty$ $,\quad T,R\in\mathbb{N},$ (6.1)
$E[0\vee(-\log\rvert X_{0}^{i}-X_{0}^{j}\rvert);X_{0}^{i},X_{0}^{j}\in S_{R}]<\infty$ $,\quad i\neq j,R\in\mathbb{N}.$ (6.2)
Let $\mathsf{X}_{t}=\sum_{i\in\mathbb{N}}\delta_{X_{t}^{i}}$ and $\mathsf{S}_{\mathrm{s}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\{{x}\})\in\{0,1\}\text{ for all }{x}\in\mathbb{R}^{d}\}$. Then
@@ 6. Non-collision in ISDEs with infinitely many particles
Let $i\neq j\in\mathbb{N}$ be fixed. Without loss of generality, we can assume $X_{0}^{i},X_{0}^{j}\in S_{R}$. We divide the case into two parts: $d\geq 3$ and $d=2$.
Suppose $d\geq 3$. Let
Let $\sum_{\langle i,j\rangle}=\sum_{(k,l)=(i,j),(j,i)}$. From (2.32), (2.33), and Itô’s formula,
$=-\sum_{\langle i,j\rangle}E\Big[\int_{0}^{t\wedge\tau_{\epsilon}}\Big(\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}},\mathfrak{b}(\mathscr{X}_{u}^{k})\Big)_{\mathbb{R}^{d}}du\Big]$
$-\frac{d-2}{2}\sum_{\langle i,j\rangle}E\Big[\int_{0}^{t\wedge\tau_{\epsilon}}\Big(\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}}\mathfrak{a}(X_{u}^{k}),\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}}\Big)_{\mathbb{R}^{d}}du\Big].$ (6.4)
@@
[…]
$\quad\quad\quad+\sum_{\langle i,j\rangle}E\Big[\int_{0}^{t\wedge\tau_{\epsilon}}\Big(\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}},\mathfrak{b}(\mathscr{X}_{u}^{k})\Big)_{\mathbb{R}^{d}}du\Big].$ (6.5)
Let $c_{6.1}$ and $c_{6.2}$ be the positive constants defined by
@@
By (6.8) and $d\geq 3$, we have $\limsup_{\epsilon\to 0}x_{\epsilon}<\infty$ and $\limsup_{\epsilon\to 0}y_{\epsilon}<\infty$. Hence by Fatou’s lemma and $\limsup_{\epsilon\to 0}x_{\epsilon}<\infty$, for each $t\geq 0$,
@@
$=\liminf_{\epsilon\to 0}x_{\epsilon}<\infty.$ (6.9)
$E[0\vee(-\log\lvert X_{\tau_{0}}^{i}-X_{\tau_{0}}^{j}\rvert);\tau_{0}>0]$
Let $\tilde{\tau}_{R,T}^{i,j}=\inf\{0\leq t\leq T;X_{t}^{i}\notin S_{R}\text{ or }X_{t}^{j}\notin S_{R}\}$. By (6.10),
$P(X_{t}^{i}=X_{t}^{j}\text{ for some }0<t<\tilde{\tau}_{R,T}^{i,j};\tau_{0}>0)=0$
for all $R,T,i\neq j\in\mathbb{N}$. This implies (6.3) for $d\geq 3$.
We next suppose $d=2$. Applying Itô’s formula to $)\log(0\vee(-\log\lvert{x}\rvert$)) instead of $\log\lvert{x}\rvert$ in (6.4) yields (6.3) for $d=2$ in a similar fashion. For the sake of completeness, we present further details.
@@
$=-\sum_{\langle i,j\rangle}E\Big[\int_{0}^{t\wedge\tau_{\epsilon}}\frac{1}{v(X_{u}^{k}-X_{u}^{l})}\Big(\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}},\mathfrak{b}(\mathscr{X}_{u}^{k})\Big)_{\mathbb{R}^{d}}du\Big]$
$-\sum_{\langle i,j\rangle}E\Big[\int_{0}^{t\wedge\tau_{\epsilon}}\frac{1}{v(X_{u}^{k}-X_{u}^{l})^{2}}\Big(\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}}\mathfrak{a}(X_{u}^{k}),\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}}\Big)_{\mathbb{R}^{d}}du\Big].$
From (2.31) and the display equation right above, we obtain
[…]
$\leq E[\log v(X_{0}^{i}-X_{0}^{j})]$
$-\sum_{\langle i,j\rangle}E\Big[\int_{0}^{t\wedge\tau_{\epsilon}}\frac{1}{v(X_{u}^{k}-X_{u}^{l})}\Big(\frac{X_{u}^{k}-X_{u}^{l}}{\lvert X_{u}^{k}-X_{u}^{l}\rvert^{2}},\mathfrak{b}(\mathscr{X}_{u}^{k})\Big)_{\mathbb{R}^{d}}du\Big].$ (6.11)
Let $x_{\epsilon},y_{\epsilon}\geq 0$ be such that
@@
From (6.13), we have $\limsup_{\epsilon\to 0}x_{\epsilon}<\infty$ and $\limsup_{\epsilon\to 0}y_{\epsilon}<\infty$. Hence, using Fatou’s lemma, we obtain that, for each $R\in\mathbb{N}$ and $t\geq 0$,
[…]
$=\liminf_{\epsilon\to 0}x_{\epsilon}<\infty.$
Proposition 6.2.
Let $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ be a weak solution of (2.33). Assume
$P\circ\mathbf{X}_{0}^{-1}\ll\mu\circ\mathfrak{l}^{-1}$ (6.14)
and (6.1). Then (6.3) holds.
@@
From (6.1), (6.14), and Fubini’s theorem,